# Detecting the tidal heating with the generic extreme-mass-ratio inspirals

### v2

@@ Abstract
The horizon of a classical black hole, functioning as a one-way membrane, plays a vital role in the dynamic evolution of binary objects, capable of absorbing fluxes entirely. Tidal heating, stemming from this phenomenon, exerts a notable influence on the production of gravitational waves (GWs). This impact can be utilized for model-independent investigations into the nature of massive objects. In this paper, assuming that the extreme-mass-ratio inspiral (EMRI) contains a stellar-mass compact object orbiting around a massive exotic compact object (ECO) with a reflective surface, we compute the GWs from the generic EMRI orbits. Using the accurate and analytic flux formulas in the black hole spacetime, we adapted these formulas in the vicinity of the ECO surface by incorporating a reflectivity parameter. Under the adiabatic approximation, we can evolve the orbital parameters and compute the EMRI waveforms. The effect of tidal heating for the spinning and non-spinning objects can be used to constrain the reflectivity of the surface at the level of $\mathcal{O}(10^{-4})$ by computing the mismatch and Fisher information matrix.
@@ I Introduction
In the framework of classical general relativity (GR), the event horizon of a black hole (BH), as a null hypersurface, can absorb all radiation near the horizon. The famous Hawking radiation stated that there would exist the non-trivial quantum effects due to the causal structure of the horizon Hawking (1976). From the viewpoint of quantum gravity, the spacetime structure near the horizon could be corrected at the Planck scale Bianchi et al. (2021); Giddings (2017); Maselli et al. (2018); Agullo et al. (2021); Bianchi et al. (2021, 2020). There is another possibility that the extremely compact objects can be formed via the phase transition Mazur and Mottola (2004, 2004). All these configuration would result in the similar objects, where the external spacetime of compact objects behaves almost the same as BHs, except for that near the horizons. These BH-like compact objects are usually called as exotic compact objects (ECOs) Cardoso and Pani (2019).
@@
So far, the nearly one hundred GW events emitted from the coalesces of compact binaries have been reported, which provided the novel ways to test GR and the chances to find the evidences beyond GR Abbott et al. (2019, 2021a, 2021b, 2021c). Nowadays one can confirm that there indeed exist the dark compact objects (DCOs) in the universe, their spacetime and dynamics can be well described by Kerr BH according to the no-hair theorem Kerr (1963); Bekenstein (1996); Chrusciel et al. (2012). The observation of binary DCOs can reach scales comparable to the sizes of BHs. The Event Horizon Telescope (EHT) also provides a new means to probe the near-horizon physical effect Akiyama et al. (2019); Gralla et al. (2019); Akiyama et al. (2022). According to null-hypothesis tests and searches for signatures of GR violations, no reliable evidence has been found in the observation data sets detected by LIGO, Virgo, and EHT. It is very likely because the current observed GW events have the relatively small signal-to-noise ratios (SNRs). The future space-based and ground-based gravitational wave detectors expected in the next decade, such as the Einstein Telescope Punturo et al. (2010), LISA Amaro-Seoane et al. (2017), and TianQin Luo et al. (2016); Gong et al. (2021), will be capable of observing various types of signals with higher SNRs originating from more massive gravitational sources. This capability will enable accurate tests of GR and the strong constraint of its alternatives in the strong field regime or near the horizon.
@@
Due to the mechanism of superradiance, a rotating BH can not only absorb radiations but also amplify radiations of lower frequency. For a more detailed introduction, please refer to the review by Brito $et~{}al$. Brito et al. (2015). This phenomenon makes BHs become the dissipative systems, like the Newtonian viscous fluid Poisson (2009); Cardoso and Pani (2013). Because of dissipation, BHs would produce the novel effects in the binaries, such as tidal heating Hartle (1973); Hughes (2001); Comeau and Poisson (2009); Poisson (2015); Landry and Poisson (2015); Chatziioannou et al. (2016); Pitre and Poisson (2023). During the inspiral of binary objects, the GW emissions are easily subjected with the surrounding tidal field, especially at the phase of plunge and merger. Tidal heating arises when GW radiation is absorbed to some extent by one object, a process that can be conceptualized as an interaction with the tidal field generated by its companion. This tidal interaction induces deformation in the object, which subsequently influences the companion’s orbit. As a result, energy and angular momentum are transferred from the object to the orbit, contributing to the overall dynamics of the system Hartle (1973); Datta et al. (2020). Assuming that the binary objects include at least one ECO, it would be possible for the tidal heating to exert an influence on the inspiraling orbital dynamics, which might become more evident for the rapidly rotating objects Datta et al. (2020). With the incorporation of tidal heating, the exploration of physical effects at the horizon scale becomes feasible through observations made by LISA Maselli et al. (2018); Johnson-Mcdaniel et al. (2020); Cardoso and Pani (2019).
@@
An EMRI consisting of a stellar-mass compact object (with mass $m\sim 1-30M_{\odot}$) inspiraling around a massive compact object (MCO) (with mass $M\sim 10^{4}-10^{7}M_{\odot}$) can emit the $m\rm Hz$ GW signals in the band of LISA, which can sever as a ideal tool to map the spacetime geometry of the central MCO Ryan (1995); Berry et al. (2019). In the realm of EMRIs, the inclusion of tidal heating significantly influences orbital evolution. This influence opens avenues for a model-independent test to discern the presence or absence of a horizon Datta and Bose (2019); Datta et al. (2020). In addition, alternative tests for determining the nature of MCOs in the context of EMRI encompass examinations of tidal deformability Pani and Maselli (2019); De Luca et al. (2023); Zi and Li (2023) and area quantization Datta and Phukon (2021). Datta $et~{}al.$ argued that the tidal heating for the circular and equatorial EMRI orbits allows to constrain the reflectivity of the ECO at the level of $\mathcal{O}(0.01\%)$ by computing the mismatch between the waveforms from Kerr BH and the ECO Datta et al. (2020). Additionally, modifying the boundary conditions near the horizon is a feasible method to probe the nature of MCOs. Recently, Maggio $et~{}al.$ evaluate the bound on reflectivity using the EMRI around Kerr-like horizonless ECOs at the level of $\mathcal{O}(10^{-6})$, where the fluxes near the surface need to be modified by imposing the ingoing and outgoing boundary conditions simultaneously Maggio et al. (2021). Using the similar configuration, Sago $et~{}al.$ found that there indeed exists the anomalous oscillating behaviors in the fluxes for the horizonless ECOs Sago and Tanaka (2021), which are expected to be a signal of the presence of a reflective surface Sago and Tanaka (2022).
@@
Note that the above analysis regarding the EMRI-based constraint on the reflectivity of ECOs only focus on the equatorial and circular orbits. However, the typical EMRI systems are the inclined and eccentric binary objects since they are birthed in the center of galaxies Berry et al. (2019). In this paper, we plan to extend the assumption in the Ref. Datta et al. (2020), considering the generic EMRI orbits and presenting the more rigorous constraint on the tidal heating parameter with the Fisher information matrix (FIM) rather than the mismatch. Firstly, assuming that external spacetime of the MCO with a reflective surface can be approximately described by Kerr BH or Schwarzschild BH, we will compute the parameterized fluxes near the surface in term of the analytic accurate flux formulas for the Kerr BH Sago and Fujita (2015) and for the Schwarzschild BH Munna et al. (2020, 2023). Then the modified fluxes for the MCO can be embedded into the FastEMRIWaveform (few) model to generate the EMRI waveform rapidly. In order to evaluate the effect of tidal heating on the EMRI orbital phase and waveform, we will calculate dephasing and mismatch between waveforms from the ECO and Kerr or Schwarzschild BH. Finally, to assess the capability of LISA of detecting the tidal heating effect, we will place the constraint on the reflectivity using the FIM method.
@@ II.1 Background of the massive compact object and geodesics
Firstly we introduce the geodesics of a timelike particle near the ECO. The orbits can be described by the Mino time $d\lambda$, which is defined via proper time $d\tau$ by $d\lambda=d\tau/\Sigma$. Then the geodesic equations of timelike particles are given by Mino (2003),
$\frac{dr}{d\lambda}$ $=$ ±$\pm\sqrt{R},$ (2)
$\frac{d\theta}{d\lambda}$ $=$ ±$\pm\sqrt{\Theta},$ (3)
@@
$T_{\theta}$ $=$ $\frac{(r^{2}+a^{2})^{2}}{\Delta}E-\frac{2MraL_{z}}{\Delta},$ (8)
$T_{r}$ $=$ $-a^{2}\sin^{2}\theta E,$ (9)
@@
where $E$, $L_{z}$ and $Q$ is the energy, angular momentum and Carter constant of the particles. Their analytic expressions and the fundamental orbital frequencies have been derived in Refs. Schmidt (2002); Fujita and Hikida (2009).
From the Eqs. (2) and (3), the orbits oscillate in the ranges $r_{\rm min}\leq r\leq r_{\rm max}$ and $\theta_{\rm min}\leq\theta\leq\theta_{\rm max}$ with $\theta_{\rm max}=\pi-\theta_{\rm min}$ and $r_{\rm min/max}=p/(1\pm e)$, where $p$ is the orbital semi-latus rectum and $e$ is the orbital eccentricity, and the orbital inclination $x_{I}$ can be obtained by $\theta_{\rm min}$ Drasco and Hughes (2006); Hughes et al. (2021), defined by $x_{I}=\pi/2-\mathrm{sgn}(L_{z})\theta_{\rm min}$.
@@ II.2 Fluxes of ECOs
In our hypothesis, the surface has an indirect influence on the orbits around the ECO and only modifies the fluxes near the surface. The EMRI trajectories are a series of geodesics of the Kerr BH, whose evolution are subjected with fluxes formulas obtained by solving Teukolsky equation Hughes et al. (2021); Isoyama et al. (2022). However, obtaining adiabatic EMRI waveforms comes with a higher computational cost, primarily due to the presence of the multivoice structure Hughes et al. (2021). The analytic post-Newtonian (PN) formulas of energy and angular fluxes have been derived up to the 4PN order in Kerr BH Sago and Fujita (2015) and to the higher order in the Schwarzschild BH Munna et al. (2020, 2023). Assuming that the orbits carry the low-velocity and smaller eccentricity, the relative errors in the analytic PN formulas can reach to the tolerable level Sago and Fujita (2015); Munna et al. (2020, 2023); Isoyama et al. (2022).
@@
where $\mathcal{C}\in(E,L_{z},Q)$ and $\dot{\mathcal{C}}$ denote the energy flux, angular momentum flux and Carter constant fluxes near the surface of the ECO. The parameter $\mathcal{R}$ denotes the reflectivity coefficient of surface Maggio et al. (2017); Mark et al. (2017). For simplicity, we assume that $\mathcal{R}$ is a constant quantity independent of the frequency. The parameter $\mathcal{R}$ vanishes for a BH, however, it can be equal to one for a fully reflective surface of a ECO. On the other hand, for the fluxes at infinity, we make the hypothesis that the fluxes at infinity for the ECO keep the same as those of the Kerr BH.
@@
$\frac{d\mathcal{C}}{dt}^{\rm orbit}=-\frac{d\mathcal{C}^{\infty}}{dt}-\frac{d\mathcal{C}^{H}}{dt}.$ (13)
@@ II.3 Orbital adiabatic evolution
$\frac{dp}{dt}$ $=$ $\frac{(1-e^{2})}{2p}\left[(1-e)\frac{dr_{a}}{dt}-(1+e)\frac{dr_{p}}{dt}\right],$ (15)
$\frac{de}{dt}$ $=$ $\frac{(1-e)^{2}}{2}\frac{dr_{a}}{dt}+\frac{(1+e)^{2}}{2}\frac{dr_{p}}{dt},$ (16)
@@
$\frac{dr_{a,p}}{dt}=J_{E_{r_{a,p}}}\frac{dE}{dt}+J_{L_{z}{{}_{r_{a,p}}}}\frac{dL_{z}}{dt}+J_{Q_{r_{a,p}}}\frac{dQ}{dt}$ (17)
@@
Here the $(E,L_{z},Q)$ and $(\dot{E},\dot{L_{z}},\dot{Q})$ are the geodesic constants and their variation rates. The analytic expressions of geodesic constants $(E,L_{z},Q)$ have been derived in Ref. Schmidt (2002); van de Meent (2020) and their change rates (13).

### v6

# Detecting the tidal heating with the generic extreme mass-ratio inspirals

@@ Abstract
The horizon of a classical black hole (BH), functioning as a one-way membrane, plays a vital role in the dynamic evolution of binary BHs, capable of absorbing fluxes entirely. Tidal heating, stemming from this phenomenon, exerts a notable influence on the production of gravitational waves (GWs). If at least one member of a binary is an exotic compact object (ECO) instead of a BH, the absorption of fluxes is expected to be incomplete and the tidal heating would be different. Thus, tidal heating can be utilized for model-independent investigations into the nature of compact object. In this paper, assuming that the extreme mass-ratio inspiral (EMRI) contains a stellar-mass compact object orbiting around a massive ECO with a reflective surface, we compute the GWs from the generic EMRI orbits. Using the accurate and analytic flux formulas in the black hole spacetime, we adapted these formulas in the vicinity of the ECO surface by incorporating a reflectivity parameter. Under the adiabatic approximation, we can evolve the orbital parameters and compute the EMRI waveforms. The effect of tidal heating for the spinning and non-spinning objects can be used to constrain the reflectivity of the surface at the level of $\mathcal{O}(10^{-6})$ by computing the mismatch and fisher information matrix.
@@ I Introduction
In the framework of classical general relativity (GR), the event horizon of a black hole (BH), as a null hypersurface, can absorb all radiation near the horizon. The famous Hawking radiation stated that there would exist the non-trivial quantum effects due to the causal structure of the horizon Hawking (1976). From the viewpoint of quantum gravity, the spacetime structure near the horizon could be corrected at the Planck scale Bianchi et al. (2021); Giddings (2017); Maselli et al. (2018); Agullo et al. (2021); Bianchi et al. (2021, 2020). There is another possibility that the extremely compact objects can be formed via the phase transition Mazur and Mottola (2004, 2004). All these configuration would result in the similar objects, where the external spacetimes of compact objects behave almost the same as BHs, except for that near the horizons. These BH-like compact objects are usually called as exotic compact objects (ECOs) Cardoso and Pani (2019).
@@
So far, the nearly one hundred GW events emitted from the coalesces of compact binaries have been reported, which provided the novel ways to test GR and the chances to find the evidences beyond GR Abbott et al. (2019a, 2021a, 2021b, 2021c). Nowadays ones can confirm that there indeed exist massive BHs in the universe Akiyama et al. (2019a, a, 2022a, b), their spacetime and dynamics can be well described by the Kerr metric according to the no-hair theorem Kerr (1963); Bekenstein (1996); Chrusciel et al. (2012). The Event Horizon Telescope (EHT) also provides a new means to probe the near-horizon physical effect Akiyama et al. (2019a); Gralla et al. (2019); Akiyama et al. (2022b). According to the null-hypothesis testings and searches for signatures of GR violations, no reliable evidence has been found in the observation data sets detected by LIGO, Virgo, and EHT Abbott et al. (2019a, b, 2021a, 2021c); Ghosh (2022); Psaltis et al. (2020); Akiyama et al. (2022b). It is very likely because the current observed GW events have the relatively small signal-to-noise ratios (SNRs). The future space-based and ground-based gravitational wave detectors expected in the next decade, such as the Einstein Telescope Punturo et al. (2010), LISA Amaro-Seoane et al. (2017), and TianQin Luo et al. (2016); Gong et al. (2021), will be capable of observing various types of signals with higher SNRs originating from more massive gravitational sources. This capability will enable accurate tests of GR and the strong constraint of its alternatives in the strong field regime or near the horizon.
@@
In Newtonian gravity, for a two-body system, orbital and rotational energy is dissipated as heat through the process of tidal friction, known as tidal heating. Similarly, for a classical black hole, the event horizon acts as a one-way membrane that can completely absorb incoming fluxes. This absorption by the BH’s horizon can also be referred to as tidal heating Hartle (1973); Hughes (2001); Comeau and Poisson (2009); Poisson (2015); Landry and Poisson (2015); Chatziioannou et al. (2016); Pitre and Poisson (2023). In addition, in Newtonian gravity, tidal heating for the rotating bodies is always accompanied by the transfer of energy and angular momentum from the bodies to their orbit. For the rotating BHs case, this phenomenon would also happen. Due to the mechanism of superradiance Brito et al. (2015), a rotating BH can not only absorb radiations but also amplify radiations of lower frequency. This then makes BHs become the dissipative systems, like the Newtonian viscous fluid Poisson (2009); Cardoso and Pani (2013). During the inspiral of a binary BH, the radiation absorbed by one BH can be understood as the tidal field generated by the companion. This tidal interaction induces deformation on the horizon of the BH, especially at the late stage of the inspiral, which subsequently influences the companion’s orbit, just like the bulge on the earth acts on motion of the moon Hughes (2001); O’Sullivan and Hughes (2016). As a result, as long as the BH rotates fast enough such that the superradiance condition is satisfied, energy and angular momentum are transferred from the BH into the orbit, contributing to the overall dynamics of the system Hartle (1973); Datta et al. (2020). Assuming at least one member of a binary is an ECO, the absorption of radiation by this ECO could be significant reduced, depending on the reflectivity of its surface. Then the tidal heating and the inspirals in the binary would be also affected, which for the rapidly rotating objects might become more evident Datta et al. (2020). Therefore, tidal heating provides a model-independent test to discern the presence or absence of a horizon. With the incorporation of tidal heating, one expects that the exploration of physical effects at the horizon scale could be feasible through observations made by LISA Maselli et al. (2018); Johnson-Mcdaniel et al. (2020); Cardoso and Pani (2019).
@@
An EMRI consisting of a stellar-mass compact object (with mass $m\sim 1-10^{2}M_{\odot}$) inspiraling around a massive compact object (MCO) (with mass $M\sim 10^{5}-10^{9}M_{\odot}$) can emit the $m\rm Hz$ GW signals in the band of LISA, which can sever as an ideal tool to map the spacetime geometry of the central MCO Ryan (1995); Berry et al. (2019). To determine the nature of the MCO in the EMRI system, especially whether the horizon exists, various approaches have been proposed. If the MCO is an ECO instead of a BH, then the nonvanishing of tidal deformability would leave an observable imprint on the EMRI waveform Pani and Maselli (2019); De Luca et al. (2023); Zi and Li (2023). However, to our knowledge these works are limited to the nonrelativistic case, which may not be accurate to capture the physics in the strong-field regime. Moreover, tidal heating provides a model-independent tool to discern the presence of horizons for the MCOs in the EMRIs Datta and Bose (2019); Datta et al. (2020). In particular, Datta $et~{}al.$ Datta et al. (2020) argued that the tidal heating for the circular and equatorial EMRI orbits allows to constrain the reflectivity of the ECO at the level of $\mathcal{O}(0.01\%)$ by computing the mismatch between the relativistic waveforms from the Kerr BH and the ECO spacetimes Datta et al. (2020). Further works were focused on modifying the boundary conditions near the surface of the ECOs. For example, Maggio $et~{}al.$ Maggio et al. (2021) evaluated the bound on reflectivity using the EMRI around Kerr-like horizonless ECOs at the level of $\mathcal{O}(10^{-6})$, where the fluxes near the surface need to be modified by imposing the ingoing and outgoing boundary conditions simultaneously. Using the similar configuration, Sago $et~{}al.$ found that there indeed exists the anomalous oscillating behaviors in the fluxes for the horizonless ECOs Sago and Tanaka (2021), which is expected to be a signal of the presence of a reflective surface Sago and Tanaka (2022).
Note that the above analyses using the tidal heating within the EMRI dynamics to constrain the reflectivity of ECOs only focus on the equatorial and circular orbits. However, the typical EMRI systems are the inclined and eccentric binary objects when they are birthed in the center of galaxies Berry et al. (2019). Therefore it is necessary and important to extend the work in Ref. Datta et al. (2020) by considering the generic EMRI orbits. Furthermore, it is also important to present the more rigorous constraint on the tidal heating parameter with the fisher information matrix (FIM) Cutler and Flanagan (1994); Vallisneri (2008) rather than the mismatch.
@@
We will assume that the MCO of the EMRI has a reflective surface instead of a horizon and the external spacetime is approximately described by the Kerr metric or Schwarzschild metric. We will compute the parameterized fluxes near the surface in term of the analytic accurate flux formulas for the Kerr BH Sago and Fujita (2015) and for the Schwarzschild BH Munna et al. (2020, 2023). Then the modified fluxes for the MCO can be embedded into a latest EMRI waveform model that can be used for data analysis, referred to as FastEMRIWaveform (FEW) Katz:2021yft to generate the EMRI waveform rapidly. In order to evaluate the effect of tidal heating on the EMRI orbital phase and waveform, we will calculate dephasing and mismatch between waveforms from the ECO and Kerr or Schwarzschild BH. Finally, to assess the capability of LISA for detecting the tidal heating effect, we will place the constraint on the reflectivity using the FIM method.
@@ II.1 Background of the massive compact object and geodesics
Firstly we introduce the geodesics of a timelike particle near the ECO. The orbits can be described by the Mino time $d\lambda$, which is defined via proper time $d\tau$ by $d\lambda=d\tau/\Sigma$. Then the geodesic equations of timelike particles are given by Thorne et al. (1973); Drasco and Hughes (2006),
$\left(\frac{dr}{d\lambda}\right)^{2}$ $=$ $R,$ (2)
$\left(\frac{d\theta}{d\lambda}\right)^{2}$ $=$ $\Theta,$ (3)
@@
$T_{r}$ $=$ $\frac{(r^{2}+a^{2})^{2}}{\Delta}E-\frac{2MraL_{z}}{\Delta},$ (8)
$T_{\theta}$ $=$ $-a^{2}\sin^{2}\theta E,$ (9)
@@
where $E$, $L_{z}$ is the energy and angular momentum the orbit per unit mass $m$, respectively, and $Q$ is the Carter constant of the orbit per unit mass square $m^{2}$. Their analytic expressions and the fundamental orbital frequencies have been derived in Refs. Schmidt (2002); Fujita and Hikida (2009).
From the Eqs. (2) and (3), the orbits oscillate in the ranges $r_{\rm min}\leq r\leq r_{\rm max}$ and $\theta_{\rm min}\leq\theta\leq\theta_{\rm max}$ with $\theta_{\rm max}=\pi-\theta_{\rm min}$ and $r_{\rm min/max}=p/(1\pm e)$, where $p$ is the orbital semi-latus rectum and $e$ is the orbital eccentricity, and the orbital inclination angle $I$ can be obtained by $\theta_{\rm min}$ Drasco and Hughes (2006); Hughes et al. (2021), defined by $\theta_{I}=\pi/2-\mathrm{sgn}(L_{z})\theta_{\rm min}$. The parameter $x_{I}=\cos\theta_{I}$ changes from $-1$ to $1$ when the orbits transform from the retrograde equatorial to the prograde equatorial.
@@ II.2 Fluxes of ECOs
In our hypothesis, the surface has an indirect influence on the orbits around the ECO and only modifies the fluxes near the surface. The EMRI trajectories are a series of geodesics of the Kerr BH, whose evolution are subjected with fluxes formulas obtained by solving Teukolsky equation Hughes et al. (2021); Isoyama et al. (2022). However, it takes a high computational cost to obtain adiabatic EMRI waveforms when computing the FIM, this is mainly due to the longtime duration and the higher harmonic element of the EMRI waveforms for the generic orbits, the detailed discussion can refer to the paper Hughes et al. (2021). The analytic post-Newtonian (PN) formulas of energy and angular fluxes have been derived up to the 4PN order in the Kerr BH Sago and Fujita (2015) and to the higher order in the Schwarzschild BH Munna et al. (2020, 2023). If both the orbital velocity and eccentricity are small, the relative errors of the analytic PN formulas can reach to the tolerable level Sago and Fujita (2015); Munna et al. (2020, 2023); Isoyama et al. (2022).
@@
where $\mathcal{C}\in\{E,L_{z},Q\}$ and $\dot{\mathcal{C}}$ denote the energy flux, angular momentum flux and Carter constant fluxes near the surface of the ECO. The parameter $\mathcal{R}$ denotes the reflectivity coefficient of surface Maggio et al. (2017); Mark et al. (2017). For simplicity, we assume that $\mathcal{R}$ is a constant quantity independent of the frequency. The parameter $\mathcal{R}$ vanishes for a BH, however, it can be equal to one for a fully reflective surface of a ECO. On the other hand, for the fluxes at infinity, we make the hypothesis that the fluxes at infinity for the ECO keep the same as those of the Kerr BH.
@@
$\frac{d\mathcal{C}}{dt}^{\rm orbit}=-\left(\frac{d\mathcal{C}^{\infty}}{dt}+\frac{d\mathcal{C}^{H}}{dt}\right).$ (13)
@@ II.3 Orbital adiabatic evolution
$\frac{de}{dt}$ $=$ $\frac{(1-e^{2})}{2p}\left[(1-e)\frac{dr_{a}}{dt}-(1+e)\frac{dr_{p}}{dt}\right],$ (15)
$\frac{dp}{dt}$ $=$ $\frac{(1-e)^{2}}{2}\frac{dr_{a}}{dt}+\frac{(1+e)^{2}}{2}\frac{dr_{p}}{dt},$ (16)
@@
$\frac{dr_{a,p}}{dt}=J_{E_{r_{a,p}}}\dot{E}+J_{L_{z}{{}_{r_{a,p}}}}\dot{L_{z}}+J_{Q_{r_{a,p}}}\frac{dQ}{dt}$ (17)
@@
Here the expressions $\{E,L_{z},Q\}$ and $\{\dot{E},\dot{L_{z}},\dot{Q}\}$ are the geodesic constants and their variation rates. The analytic expressions of geodesic constants $\{E,L_{z},Q\}$ have been derived in Ref. Schmidt (2002); van de Meent (2020) and their change rates (13).